# BarcodeBERT: Transformers for Biodiversity Analyses

### v2

@@ Abstract
In the global challenge of understanding and characterizing biodiversity, short species-specific genomic sequences known as DNA barcodes play a critical role, enabling fine-grained comparisons among organisms within the same kingdom of life. Although machine learning algorithms specifically designed for the analysis of DNA barcodes are becoming more popular, most existing methodologies rely on generic supervised training algorithms. We introduce BarcodeBERT, a family of models tailored to biodiversity analysis and trained exclusively on data from a reference library of 1.5 M invertebrate DNA barcodes. We compared the performance of BarcodeBERT on taxonomic identification tasks against a spectrum of machine learning approaches including supervised training of classical neural architectures and fine-tuning of general DNA foundation models. Our self-supervised pretraining strategies on domain-specific data outperform fine-tuned foundation models, especially in identification tasks involving lower taxa such as genera and species. We also compared BarcodeBERT with BLAST, one of the most widely used bioinformatics tools for sequence searching, and found that our method matched BLAST’s performance in species-level classification while being 55 times faster. Our analysis of masking and tokenization strategies also provides practical guidance for building customized DNA language models, emphasizing the importance of aligning model training strategies with dataset characteristics and domain knowledge. The code repository is available at https://github.com/bioscan-ml/BarcodeBERT.
@@ 1 Introduction
The task of estimating and understanding biodiversity on our planet remains a monumental challenge, as traditional methods of taxonomic analysis often struggle to keep pace with the rate of discovery and identification of new species. In this context, the search for highly expressive, short standardized genomic regions containing meaningful taxonomic information (DNA barcodes) has become prominent in biodiversity research over the past two decades (Hebert et al., 2003; Miller, 2007; Hebert et al., 2016; Srivathsan et al., 2021). Specifically, a 658-base-pair-long fragment of the Cytochrome c Oxidase Subunit I (COI) gene (Lunt et al., 1996) has emerged as the de facto DNA barcode for kingdom Animalia (Dopheide et al., 2019) and has proven effective in addressing inherent taxonomic challenges. Particularly, barcodes can be used for fast and accurate queries to categorize novel specimens into existing taxa. Furthermore, in the absence of clear species boundaries, they can be used to systematically separate specimens into groups of closely related organisms. These clusters, known as operational taxonomic units (OTUs) correspond to groups of similar specimens and can be labelled using e.g. a Barcode Index Number (BIN) (Ratnasingham & Hebert, 2007). As it is defined systematically, such a BIN system overcomes ambiguities in traditional species labelling and thus accelerates biodiversity research. Among the numerous taxonomic groups to which DNA barcoding is applicable, invertebrates, particularly arthropods, stand out as an incredibly diverse and taxonomically complex group (Basset et al., 2012), making them the focus of many methodological studies (Badirli et al., 2021, 2023; Gong et al., 2024). The diversity and taxonomic richness of this group require specialized algorithmic approaches that can capture the taxonomic structure of the data. Consequently, biodiversity researchers are increasingly turning to machine learning methods, including convolutional neural networks (CNNs) (Badirli et al., 2021) and transformer models (Gong et al., 2024), to scale taxonomic classification of arthropods and accelerate species discovery.
@@
Transformer-based models, pretrained at scale with self-supervised learning (SSL), also referred to as “foundation models,” have found applications across diverse domains thanks to their effectiveness in learning from large unlabelled datasets (Brown et al., 2020; Touvron et al., 2021). These models are often task-agnostic and can perform well on a variety of downstream tasks after fine-tuning. Despite their success in other domains, their application for taxonomic identification using DNA barcodes has not yet been extensively explored. Moreover, most DNA-based foundation models primarily target human chromosomal DNA sequences (Zhou et al., 2023; Dalla-Torre et al., 2023; Ji et al., 2021), making them suboptimal for barcode data due to domain shift between the two data types. In particular, DNA barcodes for animals stem from a specific region in mitochondrial DNA, and patterns learned from other genomic regions may be irrelevant for taxonomic classification.
@@
To summarize our contributions, we first investigate the impact of pretraining using a large and diverse DNA barcode dataset (1 million sequences, from more than 17,000 species, across 6,700 genera) on generalization to other downstream tasks. Second, we compare BarcodeBERT against several baselines such as pretrained DNA foundation models (DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023), DNABERT-S (Zhou et al., 2024), the Nucleotide Transformer NT (Dalla-Torre et al., 2023), and HyenaDNA (Nguyen et al., 2023)), a CNN baseline following the architecture introduced by Badirli et al. (2021), and the widely used alignment-based method BLAST (Altschul et al., 1990). Third, our study provides actionable insights regarding tokenization strategies, optimal masking ratios, and the importance of application-specific pretraining for DNA language models.
@@ 2 Related Work
Many machine learning approaches, particularly in the area of representation learning, have demonstrated considerable potential in biodiversity analyses as they can embed raw DNA data into an expressive lower dimensional space. Transformer-based models (Vaswani et al., 2017), known for their ability to capture complex patterns within sequential data, have shown exceptional performance in various representation learning tasks across domains, either with or without supervision (Brown et al., 2020; Touvron et al., 2021; Caron et al., 2021). These models are especially effective in learning from vast unlabelled datasets, making them ideal candidates for the analysis of genomic data, where obtaining high-quality annotations remains challenging.
There has been a growing number of self-supervised learning-based DNA language models proposed recently, most of which are based on the transformer architecture and trained using the masked language model (MLM) objective. The first foundational model in this space, DNABERT, utilizes a BERT-based transformer architecture along with $k$-mer tokenization for genome sequence prediction tasks. Following DNABERT, other models have emerged, including the Nucleotide Transformer (Dalla-Torre et al., 2023), GENA-LM (Fishman et al., 2023), and HyenaDNA (Nguyen et al., 2023). While each model varies in architectural details, tokenization methods, and training data, their reliance on SSL and the MLM objective for pretraining remains a constant. HyenaDNA is a unique entry in this space as it uses a state-space model (SSM) based on the Hyena architecture (Poli et al., 2023) and trains it for next-token prediction (a causal MLM).
@@ 3.1 Dataset
We use a reference library for Canadian invertebrates (deWaard et al., 2019) for model training and testing, comprised of 1.5 M DNA samples from the Barcode of Life Data system (BOLD) (Ratnasingham & Hebert, 2007). The dataset was further pre-processed and subdivided as described below.
@@ 3.1.1 Data pre-processing
To ensure data integrity and consistency, we performed a series of pre-processing steps over this dataset. First, empty entries were removed. Then, following standard practices (Dalla-Torre et al., 2023), IUPAC ambiguity codes (non-ACGT symbols), including alignment gaps, were uniformly replaced with the symbol N. Duplicate sequences were removed to avoid redundancy and increase the complexity of the training and pretraining tasks. Sequences with trailing N’s were truncated. Finally, sequences falling below 200 base pairs or exhibiting over 50% N characters were excluded.
@@ 3.1.2 Data partitioning
After pre-processing, 965,289 barcode sequences from 17,464 invertebrate species, across 6,712 genera were obtained. The dataset was divided into three distinct partitions for different training and evaluation purposes: (i) Seen: This partition is intended for supervised learning pipelines, particularly to evaluate the model’s ability to classify specimens from well-represented taxa. Comprised only of samples labelled to species-level, it includes 67,267 barcodes from 1,653 arthropod species, across 500 different genera, with each species represented by at least 20 and at most 50 barcodes. The partition is further split into training (70%), testing (20%), and validation (10%) subsets. (ii) Unseen: This test partition was sampled to evaluate the models in real-world conditions where specimens from underrepresented species are frequently obtained. It only contains barcodes from “rare” species with fewer than 20 barcodes in the full reference dataset. Specifically, this partition contains 4,278 barcodes from 1,826 arthropod species, none of which are present in any other partition. Moreover, this partition contains all 500 genera labels present in the Seen partition, with up to 20 barcodes sampled per genus. The label distribution shifts are shown in Figure 1, with the Seen partition reflecting the overall dataset’s distribution and the Unseen partition exhibiting a greater diversity of rare genera. (iii) Pretrain: This partition contains the remaining 893,744 barcode sequences from 14,794 invertebrate species across 6,679 genera. Note that only 35% of the sequences in this partition contain full taxonomic annotations up to the species level. See Appendix A for more details on dataset composition.
@@ 3.2 Proposed method: BarcodeBERT
A limitation of this tokenization strategy is its sensitivity to frame shifts. For example, the $k$-mer representation of the sequence GATCGA differs entirely from that of CGATCGA, even though the sequences differ by only a one-nucleotide shift. To address this issue and make our model robust to frame shifts that may occur in practice, we introduce a data augmentation step by randomly offsetting the sequence by a value ($0\leq$ offset $<k$) during pretraining to improve generalization. Before tokenization, DNA barcodes are either padded or truncated at 660 nucleotides to ensure coverage of the barcode region in the COI gene. Finally, the tokenized sequences are fed to the model and encoded into a sequence of $768$-dimensional vectors.
@@ 4.1 Experimental setup
Bayesian zero-shot learning. We selected a species-level image classification task using the INSECT dataset (Badirli et al., 2021). This is a small multimodal dataset designed for zero-shot classification of images from unseen species using DNA as auxiliary information. DNA embeddings generated by the models are paired with pre-extracted image features to classify species in a zero-shot setup. We evaluate both embeddings from pretrained and fine-tuned models on the species classification task from the INSECT dataset. Following Badirli et al. (2021), the Bayesian zero-shot learning (BSZL) framework uses image features as priors and DNA embeddings as side information. For unseen species, the $K$-nearest seen species in the DNA embedding space are used to define local priors, allowing the Bayesian model to generate posterior predictive distributions for unseen categories. To ensure a fair comparison with prior work, image features are pre-extracted using ResNet-101 (He et al., 2016). Hyperparameter tuning for the Bayesian model is performed using the same grid search space as in Badirli et al. (2021).
@@ 4.2.1 DNA-specific categorization tasks
Our evaluation, presented in Table 1, compares several models across species-level and genus-level DNA-specific categorization tasks (fine-tuning, linear probing, 1-NN probing, BIN reconstruction through ZSC). For species-level classification, we performed a BLAST search for reference and obtained a 99.7% classification accuracy after the selection of the best hit. The performance of all fine-tuned deep learning-based models is comparable with this baseline, and all transformer models outperform the CNN model as well. DNABERT-2, DNABERT-S and BarcodeBERT achieved nearly identical accuracies over 99.7%. Notably, only BarcodeBERT continues to closely match BLAST’s performance using a linear classifier, highlighting its strength in encoding meaningful features from raw data. In genus-level 1-NN probing, BarcodeBERT achieves the highest accuracy (78.5%) among the deep learning-based models, demonstrating a superior ability to generalize across taxonomic levels. BLAST, however, performs best in this task with 83.9%. This result indicates that, without fine-tuning, BarcodeBERT captures coarser taxonomic distinctions but is limited in representing the full hierarchical taxonomic structure as illustrated in Figure 3. The ZSC task provides additional insights into the model’s understanding of the hierarchical taxonomic structure. High performance in ZSC alone indicates a learned representation’s ability to finely distinguish between closely related clusters (BINs) without necessarily capturing the higher-level taxonomy. In contrast, strong 1-NN performance at higher taxonomic levels but lower ZSC accuracy suggests that the model understands the overall topology of the hierarchical taxonomic structure, even if it lacks the granularity needed for precise clustering. DNABERT and BarcodeBERT exhibit this distinction, with BarcodeBERT achieving a more balanced performance across tasks, making it the more versatile model for comprehensive DNA barcode analysis.
@@
Table 1: Classification accuracy of DNA barcode models under different SSL evaluation strategies and different efficiency metrics. The baselines are divided into three groups: alignment-based techniques, BLAST; a deep learning-based non-SSL CNN baseline; and off-the-shelf DNA foundation models pretrained on non-barcode data. These are compared against BarcodeBERT, which is specifically pretrained on DNA barcode-based datasets. For BarcodeBERT we used the best configuration of $k=4$, with 4 attention heads, and 4 layers (4-4-4). Some models supported variable stride length; for these, the numbers in parentheses are the optimal $k$-mer values that yield the best results. We also show the throughput-per-second (TPS) of the encoders, and the total duration of the classification tasks. Numbers in boldface indicate the best result across each task, and underlined indicates second place.
@@
CNN encoder 1.8 M 934 98.2 51.8 13 47.0 55 26.8
DNABERT 88.1 M 50 ($k$=6) 99.5 ($k$=4) 47.1 248 ($k$=6) 48.1 1021 79.3
DNABERT-2 118.9 M 134 99.7 87.2 101 23.5 381 38.1
DNABERT-S 117.1 M 134 99.7 93.1 101 30.6 381 62.7
HyenaDNA-tiny 1.6 M 1167 99.2 93.5 11 37.5 44 25.8
Nucleotide Transformer 55.9 M 95 99.5 65.1 140 40.1 536 22.4
BarcodeBERT (4-4-4) 29.1 M 484 99.7 99.0 27 78.5 108 73.2
@@
Two efficiency measurements are included: throughput, defined exclusively for deep-learning-based models as the number of sequences processed per second, and total runtime for classification pipelines to ensure a fair comparison with alignment-based baselines. In terms of throughput, HyenaDNA(tiny) showcases the capabilities of state space models, demonstrating high throughput with fewer parameters. However, its classification performance is lower compared to BarcodeBERT and DNABERT-2. In total run time, our results indicate that subquadratic methods like the CNN baseline and HyenaDNA perform genus-level similarity searches (1-NN probe) 13× faster than BLAST, while BarcodeBERT is 5.5× faster than BLAST and outperforms other transformer models at this task. For species-level classification pipelines that do not include the computation of the training embeddings, transformer-based models demonstrate clear advantages over traditional baselines in terms of running time. Notably, BarcodeBERT, with a moderate parameter count, matches BLAST’s high classification accuracy ($99.7\%$) with a 55× faster running time, thus providing a well-rounded option for large-scale DNA barcode applications. All efficiency experiments were conducted using an Intel(R) Xeon(R) CPU @ 2.20GHz processor and a Quadro RTX 6000 GPU.

### v3

@@ Abstract
In the global challenge of understanding and characterizing biodiversity, short species-specific genomic sequences known as DNA barcodes play a critical role, enabling fine-grained comparisons among organisms within the same kingdom of life. Although machine learning algorithms specifically designed for the analysis of DNA barcodes are becoming more popular, most existing methodologies rely on generic supervised training algorithms. We introduce BarcodeBERT, a family of models tailored to biodiversity analysis and trained exclusively on data from a reference library of 1.5 M invertebrate DNA barcodes. We compared the performance of BarcodeBERT on taxonomic identification tasks against a spectrum of machine learning approaches, including supervised training of classical neural architectures and fine-tuning of general DNA foundation models. Our self-supervised pretraining strategies on domain-specific data outperform fine-tuned foundation models, especially in identification tasks involving lower taxa such as genera and species. We also compared BarcodeBERT with BLAST, one of the most widely used bioinformatics tools for sequence searching, and found that our method matched BLAST’s performance in species-level classification while being 55 times faster. Our analysis of masking and tokenization strategies also provides practical guidance for building customized DNA language models, emphasizing the importance of aligning model training strategies with dataset characteristics and domain knowledge. The code repository is available at https://github.com/bioscan-ml/BarcodeBERT.
@@ 1 Introduction
The task of estimating and understanding biodiversity on our planet remains a monumental challenge, as traditional methods of taxonomic analysis often struggle to keep pace with the rate of discovery and identification of new species. In this context, the search for highly expressive, short standardized genomic regions containing meaningful taxonomic information (DNA barcodes) has become prominent in biodiversity research over the past two decades (Hebert et al., 2003; Miller, 2007; Hebert et al., 2016; Srivathsan et al., 2021). Specifically, a 658-base-pair-long fragment of the Cytochrome c Oxidase Subunit I (COI) gene (Lunt et al., 1996) has emerged as the de facto DNA barcode for kingdom Animalia (Dopheide et al., 2019) and has proven effective in addressing inherent taxonomic challenges. Particularly, barcodes can be used for fast and accurate queries to categorize novel specimens into existing taxa. Furthermore, in the absence of clear species boundaries, they can be used to systematically separate specimens into groups of closely related organisms. These clusters, known as operational taxonomic units (OTUs), correspond to groups of similar specimens and can be labelled using e.g. a Barcode Index Number (BIN) (Ratnasingham & Hebert, 2007). As it is defined systematically, such a BIN system overcomes ambiguities in traditional species labelling and thus accelerates biodiversity research. Among the numerous taxonomic groups to which DNA barcoding is applicable, invertebrates, particularly arthropods, stand out as an incredibly diverse and taxonomically complex group (Basset et al., 2012), making them the focus of many methodological studies (Badirli et al., 2021, 2023; Gong et al., 2025). The diversity and taxonomic richness of this group require specialized algorithmic approaches that can capture the taxonomic structure of the data. Consequently, biodiversity researchers are increasingly turning to machine learning methods, including convolutional neural networks (CNNs) (Badirli et al., 2021) and transformer models (Gong et al., 2025), to scale taxonomic classification of arthropods and accelerate species discovery.
@@
Transformer-based models, pretrained at scale with self-supervised learning (SSL), also referred to as “foundation models,” have found applications across diverse domains thanks to their effectiveness in learning from large unlabelled datasets (Brown et al., 2020; Touvron et al., 2021). Such models are often task-agnostic and can perform well on a variety of downstream tasks after fine-tuning. Despite their success in other domains, their application for taxonomic identification using DNA barcodes has not yet been extensively explored. Moreover, most DNA-based foundation models primarily target human chromosomal DNA sequences (Zhou et al., 2023; Dalla-Torre et al., 2023; Ji et al., 2021), making them suboptimal for barcode data due to domain-shift between these data types. Though short, barcodes encode rich taxonomic information across over 265,000 animal species in the Barcode of Life Data System (BOLD).
@@
To summarize our contributions, we first investigate the impact of pretraining using a large and diverse DNA barcode dataset (1 million sequences, from more than 17,000 species, across 6,700 genera) on generalization to other downstream tasks. Second, we compare BarcodeBERT against several baselines such as pretrained DNA foundation models (DNABERT (Ji et al., 2021), DNABERT-2 (Zhou et al., 2023), DNABERT-S(Zhou et al., 2024), the Nucleotide Transformer NT (Dalla-Torre et al., 2023), and HyenaDNA (Nguyen et al., 2023), a CNN baseline following the architecture introduced by (Badirli et al., 2021), and the widely used alignment-based method BLAST (Altschul et al., 1990). Third, our study provides actionable insights regarding tokenization strategies, optimal masking ratios, and the importance of application-specific pretraining for DNA language models.
@@ 2 Related Work
Many machine learning approaches, particularly in representation learning, have demonstrated considerable potential in biodiversity analyses as they can embed raw DNA data into an expressive lower-dimensional space. Transformer-based models (Vaswani et al., 2017) capture complex patterns within sequential data and have shown exceptional performance in various representation learning tasks across domains, either with or without supervision (Brown et al., 2020; Touvron et al., 2021; Caron et al., 2021). These models are especially effective in learning from vast unlabelled datasets, making them ideal candidates for the analysis of genomic data, where obtaining high-quality annotations remains challenging.
There has been a growing number of self-supervised learning-based DNA language models proposed recently, most of which are based on the transformer architecture and trained using the masked language model (MLM) objective. The first foundational model in this space, DNABERT, utilizes a BERT-based transformer architecture along with $k$-mer tokenization for genome sequence prediction tasks. Following DNABERT, other models have emerged, including the Nucleotide Transformer (Dalla-Torre et al., 2023), GENA-LM (Fishman et al., 2025), and HyenaDNA (Nguyen et al., 2023). While each model varies in architectural details, tokenization methods, and training data, their reliance on SSL and the MLM objective for pretraining remains a constant. HyenaDNA is a unique entry in this space as it uses a state-space model (SSM) based on the Hyena architecture (Poli et al., 2023) and trains it for next-token prediction (a causal MLM).
@@ 3.1 Dataset
We use a reference library for Canadian invertebrates (deWaard et al., 2019) for model training and testing, comprised of 1.5 M DNA samples from BOLD (Ratnasingham & Hebert, 2007). The dataset was further pre-processed and subdivided as described below.
@@ 3.1.1 Data pre-processing
To ensure data integrity and consistency, we performed a series of pre-processing steps over this dataset. First, empty entries were removed. Then, following standard practices (Dalla-Torre et al., 2023), IUPAC ambiguity codes (non-ACGT symbols), including alignment gaps, were uniformly replaced with the symbol N. Duplicate sequences were removed to avoid redundancy and increase the complexity of the training and pretraining tasks. Sequences with trailing N’s were truncated. Finally, sequences falling below 200 base pairs or exhibiting over 50% N characters were excluded. Our preprocessed dataset is available to download at https://huggingface.co/datasets/bioscan-ml/CanadianInvertebrates-ML.
@@ 3.1.2 Data partitioning
After pre-processing, 965,289 barcode sequences from 17,464 invertebrate species, across 6,712 genera were obtained. The dataset was divided into three distinct partitions for different training and evaluation purposes: (i) Seen: This partition is intended for supervised learning pipelines, particularly to evaluate the model’s ability to classify specimens from well-represented taxa. Comprised only of samples labelled to species-level, it includes 67,267 barcodes from 1,653 arthropod species, across 500 different genera, with each species represented by at least 20 and at most 50 barcodes. The partition is further split into training (70%), testing (20%), and validation (10%) subsets. (ii) Unseen: This test partition was sampled to evaluate the models in real-world conditions where specimens from underrepresented species are frequently obtained. It only contains barcodes from “rare” species with fewer than 20 barcodes in the full reference dataset. Specifically, this partition contains 4,278 barcodes from 1,826 arthropod species, none of which are present in any other partition. Moreover, this partition contains all 500 genera labels present in the Seen partition, with up to 20 barcodes sampled per genus. The label distribution shifts are shown in Figure 1, with the Seen partition reflecting the overall dataset’s distribution and the Unseen partition exhibiting a greater diversity of rare genera. (iii) Pretrain: This partition contains the remaining 893,744 barcode sequences from 14,794 invertebrate species across 6,679 genera. Note that only 35% of the sequences in this partition contain full taxonomic annotations up to the species level. See Supplement A for more details on dataset composition.
@@ 3.2 Proposed method: BarcodeBERT
A limitation of this tokenization strategy is its sensitivity to frame shifts. For example, the $k$-mer representation of the sequence GATCGA differs entirely from that of CGATCGA, even though the sequences differ by only a one-nucleotide shift. To address this issue and make our model robust to frame shifts that may occur in practice, we introduce a data augmentation step by randomly offsetting the sequence by a value ($0\leq$ offset $<k$) during pretraining to improve generalization. Before tokenization, DNA barcodes are either padded or truncated at 660 nucleotides to ensure coverage of the barcode region in the COI gene (Elbrecht et al., 2019). Finally, the tokenized sequences are fed to the model and encoded into a sequence of $768$-dimensional vectors.
@@ 4.1 Experimental setup
Bayesian zero-shot learning. We selected a species-level image classification task using the INSECT dataset (Badirli et al., 2021). This is a small multimodal dataset designed for zero-shot classification of images from unseen species using DNA as auxiliary information. DNA embeddings generated by the models are paired with pre-extracted image features to classify species in a zero-shot setup. We evaluate both embeddings from pretrained and fine-tuned models on the species classification task from the INSECT dataset. Following (Badirli et al., 2021), the Bayesian zero-shot learning (BSZL) framework uses image features as priors and DNA embeddings as side information. For unseen species, the $K$-nearest seen species in the DNA embedding space are used to define local priors, allowing the Bayesian model to generate posterior predictive distributions for unseen categories. To ensure a fair comparison with prior work, image features are pre-extracted using ResNet-101 (He et al., 2016). Hyperparameter tuning for the Bayesian model is performed using the same grid search space as in (Badirli et al., 2021).
@@ 4.2.1 DNA-specific categorization tasks
Our evaluation (Table 1) compares several models across species-level and genus-level DNA-specific categorization tasks (fine-tuning, linear probing, 1-NN probing, BIN reconstruction through ZSC). For species-level classification, a BLAST search yielded 99.7% accuracy based on the best hit. The performance of all fine-tuned deep learning-based models is comparable to this baseline, and all transformer models outperform the CNN model as well. DNABERT-2, DNABERT-S and BarcodeBERT all reached over 99.7% accuracy. Notably, only BarcodeBERT continues to closely match BLAST’s performance using a linear classifier, highlighting its strength in encoding meaningful features from raw data. In genus-level 1-NN probing, BarcodeBERT achieves the highest accuracy (78.5%) among the deep learning-based models—more than double that of the same architecture without pretraining—demonstrating that the pretrained model produces richer embeddings, making BarcodeBERT far more robust for similarity searches of sequences from novel taxa. BLAST, however, performs best in this task (83.9%). This indicates that without fine-tuning, BarcodeBERT captures coarser taxonomic distinctions but is limited in representing the full hierarchical taxonomic structure as shown in Figure 3. In addition to accuracy, we computed the F1-score for each model across these tasks, and observed similar trends (see Supplement D.8).
The ZSC task provides additional insights into the model’s understanding of the hierarchical taxonomic structure. High performance in ZSC alone indicates a learned representation’s ability to finely distinguish between closely related clusters (BINs) without necessarily capturing the higher-level taxonomy. In contrast, strong 1-NN performance at higher taxonomic levels but lower ZSC accuracy suggests that the model understands the overall topology of the hierarchical taxonomic structure, even if it lacks the granularity needed for precise clustering. DNABERT and BarcodeBERT exhibit this distinction, with BarcodeBERT achieving a more balanced performance across tasks, making it the more versatile model for comprehensive DNA barcode analysis.
@@
Table 1: Classification accuracy of DNA barcode models under different SSL evaluation strategies and different efficiency metrics. The baselines are divided into three groups: alignment-based techniques, BLAST; a deep learning-based non-SSL CNN baseline; and off-the-shelf DNA foundation models pretrained on non-barcode data. These are compared against BarcodeBERT, which is specifically pretrained on DNA barcode-based datasets. For BarcodeBERT we used the best configuration of $k=4$, with 4 attention heads, and 4 layers. The DNABERT model supported variable stride length, and we show in parentheses the optimal $k$-mer value that yielded the best results. As an ablation, we also indicate the performance of the BarcodeBERT architecture when trained fully-supervised, without the self-supervised pretraining stage. Additionally, we show the throughput-per-second (TPS) of the encoders, and the total duration of the classification tasks. Numbers in boldface indicate the best result across each task, and underlined indicates second place.
@@
CNN encoder 1.8 M 934 98.2 N/A 13 56.0 55 N/A
DNABERT ($k\!=\!6$) 88.1 M 50 99.5 98.1 248 48.1 1021 81.5
DNABERT-2 118.9 M 134 99.7 95.7 101 23.5 381 51.0
DNABERT-S 117.1 M 134 99.7 96.5 101 30.6 381 62.8
HyenaDNA-tiny (d256) 1.6 M 1167 99.4 93.5 11 50.6 76 52.8
Nucleotide Transformer 55.9 M 95 99.5 96.2 140 40.1 536 28.5
BarcodeBERT w/o pretraining 29.1 M 484 99.5 N/A 27 37.9 108 N/A
BarcodeBERT (4-4-4) 29.1 M 484 99.7 99.2 27 78.5 108 79.9
@@
Two efficiency measurements are included: throughput, defined exclusively for deep-learning-based models as the number of sequences processed per second, and total runtime for classification pipelines to ensure a fair comparison with alignment-based baselines. In terms of throughput, HyenaDNA(tiny) showcases the capabilities of state space models, demonstrating high throughput with fewer parameters. However, its classification performance is lower compared to BarcodeBERT and DNABERT-2. In total run time, our results indicate that subquadratic methods like the CNN baseline and HyenaDNA perform genus-level similarity searches (1-NN probe) 13× faster than BLAST, while BarcodeBERT is 5.6 × faster than BLAST and outperforms other transformer models at this task. For species-level classification pipelines that do not include the computation of the training embeddings, transformer-based models demonstrate clear advantages over traditional baselines in terms of running time. Notably, BarcodeBERT, with a moderate parameter count, matches BLAST’s high classification accuracy ($99.7\%$) with a 55× faster running time, thus providing a well-rounded option for large-scale DNA barcode applications. All efficiency experiments were conducted using an Intel(R) Xeon(R) CPU @ 2.20GHz processor and a Quadro RTX 6000 GPU. Additional details on resource consumption can be found in Supplement C.